Joselyne00:

ABC has side lengths 18,24, and 30. Do the side lengths form a pythagorean triple?

Mercury:

in a pythagorean triple: a^2 + b^2 = c^3, where c is the hypotenuse (biggest side in a right triangle) and a and b are the legs applying this logic to your problem: does 18^2 + 24^2 = 30^2? if yes ---> pythagorean triple, if no ---> not a pythagorean triple

jhonyy9:

@Mercury there is c^2 not c^3 sorry just a little attention .... a^2 +b^2 = c^2 you wrote c^3

Mercury:

ah yes, thank you a^2 + b^2 = c^2 not c^3

jhonyy9:

np

jhonyy9:

@Joselyne00 do you understand ?

jhonyy9:

pythagorean triplet are 3,4,5 for example because 3^2 +4^2 = 5^2 yes ? 9 +16 = 25 do you understand it now ?

jhonyy9:

make the calcule 18^2 = ? 24^2 = ? 30 ^2 = ?